Two separate containers are filled with different gases. If these gases are

allowed to mix, obtain an expression for entropy of mixing per mole of the

mixture.


1
Expert's answer
2021-03-14T19:15:31-0400

If n(gas A)=n(gas B), T(gas A)=T(gas B), then

"\Delta S = S_1-S_2=2nRln2"